# Challenge LLMs to Reason About Reasoning: A Benchmark to Unveil Cognitive Depth in LLMs

### 28 Dec 2023 v1

@@ 4 Experiments
For each question-answer pair annotated, the evaluated model are supposed to decide the correctness of the solution and report the error step and error reason if any. The solution correctness and first error step is scored automatically based on the manual annotation result. Only when the evaluated model correctly identified the incorrect solution and error step will its error reason be further examined manually.
@@ 4.1 Commercial LLMs Evaluation
Among the closed-source commercial LLMs, ChatGPT and Claude series are the most well-known ones that have been through comprehensive instruction tuning (ouyang2022training; ANTHROPIC2023CLAUDE) and careful alignment process (OpenAI2023GPT4TR; Bai2022TrainingAH; Ganguli2022RedTL). For this study, we specifically evaluated GPT3-5-turbo-0613, Claude2.0, GPT4-0613. Given that these models have demonstrated over 80 percent accuracy in single-pass settings (luo2023wizardmath), our interest lies in assessing their performance under zero-shot conditions in the DiagGSM8K benchmark.
@@
In line with findings from Orca-2 (Mitra2023Orca2T) and other researchers, we recognized that task performance is significantly influenced by system-instruction design. Thus, we conducted empirical experiments with various instruction templates, selecting the most effective one (see Figure-LABEL:fig:PromptsApis) from our validation set for use across all models. The key insight here is that even though Chain-of-thought helps improving the diagnostic reasoning abilities, the expert tailored diagnosing procedure plays a crucial role. For instance, when models were prompted to solve the questions before contrasting their solutions with the provided answers, we observed a bias towards the provided solutions, adversely affecting diagnostic accuracy. Another key aspect of our experimental setup involves the sampling temperature. Unlike the greedy sampling method used in the aforementioned single-pass experiments, we found that greedy decoding substantially diminishes generation quality and diagnostic performance in DiagGSM8K. Consequently, after several iterations of testing on our validation set, we set the sampling temperature to 0.5 for all models.
@@
Table-2 presents the evaluation results for the three selected models. It’s evident that GPT3-5 trails behind both Claude2 and GPT4 across all metrics. For instance, in binary correctness prediction, GPT3-5 achieves a score of only 56.7%, while Claude2 and GPT4 score 65.6% and 78.6%, respectively. Figure-4 offers a visual representation of the incorrect solutions within DiagGSM8K, elucidating the progressively complex nature of the tasks – identifying incorrect solutions, locating the first error step, and articulating the error’s rationale. The graph reveals a downward trend in accuracy for these successive tasks. Notably, GPT3-5 shows a mere 4.3% success rate in identifying the first error step and explaining the reason behind it.
@@
Further scrutiny of Type I (False Positive Rate, FPR) and Type II (False Negative Rate, FNR) errors made by the models, as illustrated in Figure-5, uncovers distinct tendencies in each model’s performance. Claude2 exhibits a relatively balanced distribution between Type I and Type II errors. In contrast, GPT4 is prone to committing Type II errors at a rate approximately 2.5 times higher than Type I errors, indicating a tendency to incorrectly label correct solutions as incorrect. Conversely, GPT3-5 displays an opposite trend to GPT4, with a high false positive rate of 63.4%. This suggests that GPT3-5 is more inclined to uncritically accept the given solutions, irrespective of their actual correctness, as compared to the other two models. Possible interpretation could be made by observing the low accuracy of GPT3-5 in determining the error reasons. The overall uncertainty towards a solution combined with the insufficient calibration of RLHF potentially lead to the sycophancy behavior (Sharma2023TowardsUS).
@@
Despite the fact that all three evaluated models exhibit relatively similar pass@1 rates on the GSM8K dataset (e.g., GPT4 surpassing Claude2 by only 4 percent in accuracy as per luo2023wizardmath), the performance disparity becomes stark in the context of our DiagGSM8K benchmark. GPT4 demonstrates a significantly higher proficiency, being approximately 10 times more effective than GPT3-5 and 4 times more so than Claude2 in accurately diagnosing incorrect solutions from the same dataset. This pronounced variance in evaluation metrics highlights the limitations inherent in the current design of benchmark datasets, reinforcing the argument presented in Section-1 for the necessity of a new paradigm to comprehensively evaluate the reasoning abilities of contemporary LLMs.
@@
Model Eval Method Accuracy TPR TNR Step Step+Reason
Claude2 0-shot 1968/3000 962/1427 1006/1573 311/1573 173/1573
GPT3-5 0-shot 1701/3000 1125/1427 576/1573 159/1573 68/1573
GPT4 0-shot 2359/3000 985/1427 1374/1573 784/1573 644/1573
WizardMath-70B 3-shot 1187/3000 1176/1427 11/1573 4/1573 1/1573
Mammoth-70B 3-shot 1451/3000 1410/1427 41/1573 4/1573 1/1573
MetaMath-70B 3-shot 1471/3000 1305/1427 166/1573 22/1573 6/1573
llama2-70B-diag 0-shot 1609/3000 453/1427 1156/1573 323/1573 99/1573
Table 2: Evaluation Results of Models on DiagGSM8K: This table presents a detailed breakdown of each model’s performance, including True Positive Rate (TPR) and True Negative Rate (TNR). The ’Step’ column represents the number of instances where models accurately identified an incorrect solution and pinpointed the first error step. ’Step+Reason’ indicates instances where models not only located the first error step in incorrect solutions but also correctly explained the error’s rationale.
Figure 4: Performance Visualization on Incorrect DiagGSM8K Solutions: This graph depicts the performance of Claude2, GPT3-5, and GPT4 in identifying incorrect solutions in DiagGSM8K. ’TNR’ denotes the True Negative Rate. ’TN+Step’ refers to the ratio where models correctly identified an incorrect solution and located the first error step. ’TN+Step+Reason’ represents the frequency of models correctly determining solution correctness, identifying the first error step, and providing an accurate error reason.
Figure 5: Error Rate Analysis of Closed-Source Models on DiagGSM8K: This figure illustrates the True Positive Rate (TPR), True Negative Rate (TNR), False Positive Rate (FPR), and False Negative Rate (FNR) of closed-source models evaluated on DiagGSM8K, offering insights into their accuracy in various aspects of solution evaluation.
@@ 4.2 Open-sourced LLMs Evaluation
As mentioned briefly in Section 1 and 2, recent works of fine-tuning language models with program of thought, different reasoning types, and different difficulties have all proven effective in enhancing GSM8K benchmark results. However, just like GPT3-5 has similar pass@1 rate with GPT4 in GSM8K test set, will these open-sourced models specialized in math reasoning performs differently in the DiagGSM8K?
To address this query, we selected several state-of-the-art open-source models fine-tuned on the 70B Llama architecture (touvron2023llama). Namely, they are WizardMath-70B(luo2023wizardmath), Mammonth-70B(yue2023mammoth) and MetaMath-70B(yu2023metamath). WizardMath-70B underwent training on augmented problems of varying difficulty levels and was further refined through a process-oriented reinforcement learning procedure. Mammoth-70B received training on a diverse range of STEM-related problems, including those from the GSM8K dataset, which were augmented with GPT4-generated code solutions. MetaMath-70B’s training encompassed a substantial volume of augmented data, including rephrased and backward-transformed problems from the GSM8K set. Given that these models are specialized and not fine-tuned for general instructions, we employed three-shot in-context learning examples during inference to guide the models in adhering to the desired format and reasoning logic.
As demonstrated in Table-2, despite having exposure to similar training data, all three models failed miserably in the DiagGSM8K benchmark. This outcome highlights their lack of generalization capabilities when faced with problems similar to their training data but presented in different formats. Among these models, MetaMath displayed the most commendable performance, which is not entirely unexpected. Over half of its training dataset (approximately 240k instances) is derived directly from the GSM8K dataset, including a balanced mix of answer augmentations, question rephrasings, and backward transformations of questions. Despite having a training dataset with question types akin to those in DiagGSM8K (e.g., original and POT questions), the Mammoth model failed to demonstrate a deeper meta-understanding of the data, correctly answering only one out of 1573 incorrect solutions. WizardMath, trained using the Proximal Policy Optimization algorithm (Schulman2017ProximalPO) to optimize the joint reward of instruction adherence and solution process, also underperformed on our benchmark.
@@ 4.3 In Domain Finetuning
For our base model, we utilized Llama-2-70B-base, aligning with the approach of other open-source SOTA math models. We merged the GSM8K training set with the GPT-4 generated diagnostic data that composed of 5k incorrect solutions and 4k correct solutions. For fine-tuning, we adopted the Qlora method (Dettmers2023QLoRAEF), maintaining the same hyperparameters as used for MetaMath-70B. The outcomes of this approach are detailed in the last row of Table-2. Remarkably, despite the zero-shot setting and a considerably smaller training set, the fine-tuned Llama2 model outperformed all open-source models and even surpassed GPT3-5. Notably, the fine-tuned Llama2 model demonstrated a tendency distinct from that of GPT3-5 and other open-source models; it was less inclined to accept solutions but to over reject solutions regardless of the correctness. As depicted in Figure-6, of the 99 questions where the model accurately predicted both correctness and the first error step, a significant portion involved questions with POT and backward reasoning types. This is particularly noteworthy given that the model was trained exclusively on original questions.
@@
Figure 6: Problem types of incorrect solutions that llama2-70B-diag models has successfully found error step and error reason. Note the training set only includes solutions from original problem.
@@
Caution is necessary when interpreting the outcomes of in-domain fine-tuning. Although the fine-tuned model achieved results comparable to GPT3-5 post-fine-tuning, it’s important to note that the overall number of correct diagnoses for incorrect solutions remains relatively low. This underscores the challenging nature of our DiagGSM8K benchmark, where effective diagnosis across diverse solution spaces requires a comprehensive understanding of the problem. Consequently, simple fine-tuning strategies may not yield substantial improvements in performance. Another significant observation, as indicated in Table-3, is that despite outperforming GPT3-5 in DiagGSM8K, the model’s accuracy on the GSM8K test-set still trails behind that of GPT3-5. This discrepancy could be partially attributed to a lower True Positive Rate in correct solutions and a reduced efficacy in articulating error reasons for incorrect solutions. The fine-tuning process may have enabled the model to replicate diagnostic behaviors, but this approach does not seem to enhance the model’s fundamental comprehension of mathematical reasoning. Another interpretation worth considering is that language models, to a large extent, continue to function as ’stochastic parrots’ (Bender2021OnTD), lacking genuine understanding of the content they generate. It would be intriguing to explore how scaling up the diagnostic data or employing more sophisticated training methodologies might alter these outcomes. We will delve further into this discussion in the following section.
Models GSM8k Diag-SR
GPT4 92.0% 644/1573
Claude2 88.0% 173/1573
GPT3-5 80.8% 68/1573
WizardMath-70B 81.6% 1/1573
Mammoth-70B 76.7% 1/1573
MetaMath-70B 82.3% 6/1573
@@
Table 3: Comparison of the performances of SOTA models on GSM8K and DiagGSM8K. The GSM8K results are retrieved from corresponding paper. Diag-SR stands for the correct ratio that models successfully found the erorr step and error reason for the incorrect solutions. llama2-70B-GSM8K are llama2-70B model finetuned on the training set of GSM8k only. llama2-70B-diag are llama2-70B model finetuned on the GSM8k training set and its diagnosis augmentation by GPT4.
5 Discussion and Limitations
Exploring the ’Reason About Reasoning’ Paradigm: A Robust Indicator of Reasoning Capability and Conceptual Mastery?
In fine-tuning models for any given task, the goal often extends beyond achieving a good fit within the training distribution; it also encompasses attaining substantial generalization to unseen instances. This is particularly crucial in complex reasoning tasks, where the focus transcends mere performance metrics on a test set. What becomes equally important is the validity and logic of the reasoning process employed by the evaluated model. Consider this analogy: Understanding the radix/base system in numerals should make operations like addition in base-10 or base-9 similarly challenging. Likewise, for individuals adept at using scratchpads for calculations, the complexity of multiplying two-digit versus three-digit numbers shouldn’t differ drastically. Yet, instances of GPT-4 failing to generalize in such scenarios (Wu2023ReasoningOR; Dziri2023FaithAF) point towards a deficiency in the underlying mastery of these concepts, a flaw that mere accuracy metrics in conventional benchmarks may fail to reveal. This is exacerbated by the current LLMs’ ability to produce well-structured chain of thought responses, giving an illusion of proficiency in various reasoning tasks despite potential gaps in understanding.
This observation underscores a key issue: an intelligent agent that aligns epistemically with human reasoning is expected to base its logic on abstracted rules and principles and apply these deductively across varied contexts. Therefore, the ’reason about reasoning’ paradigm emerges as a vital meta-evaluative tool. It allows us to delve deeper into the reasoning process itself, probing the model’s ability to apply learned concepts flexibly and logically across different scenarios, and unmasking superficial understandings that might be concealed by high performance in traditional test settings.
Limitations of the ’Reason About Reasoning’ Evaluation Paradigm and DiagGSM8K Dataset

### 20 Jan 2024 v2

@@ 4.1 Commercial LLMs Evaluation
Among the closed-source commercial LLMs, ChatGPT and Claude series are the most well-known ones that have been through comprehensive instruction tuning (Ouyang et al., 2022; Anthropic, 2023) and careful alignment process (OpenAI, 2023; Bai et al., 2022; Ganguli et al., 2022). For this study, we specifically evaluated GPT3-5-turbo-0613, Claude2.0, GPT4-0613. Given that these models have demonstrated over 80 percent accuracy in single-pass settings (Luo et al., 2023), our interest lies in assessing their performance under zero-shot conditions in the MR-GSM8k benchmark.
@@
In line with findings from Orca-2 (Mitra et al., 2023) and other researchers, we recognized that task performance is significantly influenced by system-instruction design. Thus, we conducted empirical experiments with various instruction templates, selecting the most effective one (see Figure-LABEL:fig:PromptsApis) from our validation set for use across all models. The key insight here is that even though Chain-of-thought helps improving the diagnostic reasoning abilities, the expert tailored diagnosing procedure plays a crucial role. For instance, when models were prompted to solve the questions before contrasting their solutions with the provided answers, we observed a bias towards the provided solutions, adversely affecting diagnostic accuracy. Another key aspect of our experimental setup involves the sampling temperature. Unlike the greedy sampling method used in the aforementioned single-pass experiments, we found that greedy decoding substantially diminishes generation quality and diagnostic performance in MR-GSM8k. Consequently, after several iterations of testing on our validation set, we set the sampling temperature to 0.5 for all models.
@@
Table-2 presents the evaluation results for the three selected models. It’s evident that GPT3-5 trails behind both Claude2 and GPT4 across all metrics. For instance, in binary correctness prediction, GPT3-5 achieves an MCC score of only 0.198, while Claude2 and GPT4 score 0.345 and 0.614, respectively. Figure-4 offers a visual representation of the incorrect solutions within MR-GSM8k, elucidating the progressively complex nature of the tasks – identifying incorrect solutions, locating the first error step, and articulating the error’s rationale. The graph reveals a downward trend in accuracy for these successive tasks. Notably, GPT3-5 shows a mere 4.64% success rate in identifying the first error step and explaining the reason behind it.
@@
Further scrutiny of Type I (False Positive Rate, FPR) and Type II (False Negative Rate, FNR) errors made by the models, as illustrated in Figure-5, uncovers distinct tendencies in each model’s performance. Claude2 exhibits an impressively balanced distribution between Type I and Type II errors. In contrast, GPT4 is prone to committing Type II errors at a rate approximately three times higher than Type I errors, indicating a tendency to incorrectly label correct solutions as incorrect. Conversely, GPT3-5 displays an opposite trend to GPT4, with a high false positive rate of 60.52%. This suggests that GPT3-5 is more inclined to uncritically accept the given solutions, irrespective of their actual correctness, as compared to the other two models. Possible interpretation could be made by observing the low accuracy of GPT3-5 in determining the error reasons. The overall uncertainty towards a solution combined with the potentially insufficient calibration of RLHF lead to the sycophancy behavior (Sharma et al., 2023).
@@
Despite the fact that all three evaluated models exhibit relatively similar pass@1 rates on the GSM8K dataset (e.g., GPT4 surpassing Claude2 by only 4 percent in accuracy as per Luo et al., 2023), the performance disparity becomes stark in the context of our MR-GSM8k benchmark. GPT4 demonstrates a significantly higher proficiency, being approximately 5 times more effective than GPT3-5 and 2.5 times more so than Claude2 measured by MR-Score. This pronounced variance in evaluation metrics highlights the limitations inherent in the current design of benchmark datasets, reinforcing the argument presented in Section-1 for the necessity of a new paradigm to comprehensively evaluate the reasoning abilities of contemporary LLMs.
@@
Table 2: Evaluation Results of Models on MR-GSM8k: This table presents a detailed breakdown of each model’s performance, including True Positive Rate (TPR), True Negative Rate (TNR) and Matthews Correlation Coefficient. The ’Step’ column represents the accuracy of correctly identifying an incorrect solution and pinpointing the first error step. ’S+R/M’ column showcased the accuracy of not only locating the first error step in incorrect solutions but also correctly explaining the error’s rationale. The overall MR-Score/M is a normalized metric ranges from 0 to 1 and calculated based on formula described in Section-3. The M here means the error reason is labelled manually and MR-Score is calculated based on the results of that.
Model Eval Method TPR TNR MCC Step S+R/M MR-Score/M
Claude2 0-shot 67.41% 67.13% 0.345 21.04% 11.76% 0.191
GPT3-5 0-shot 78.84% 39.48% 0.198 11.38% 4.64% 0.097
GPT4 0-shot 69.03% 90.59% 0.614 52.32% 43.04% 0.495
WizardMath-70B 3-shot 82.41% 2.73% -0.250 0.38% 0.06% 0.001
Mammoth-70B 3-shot 98.81% 2.73% 0.055 0.25% 0.06% 0.012
MetaMath-70B 3-shot 91.45% 10.55% 0.034 1.40% 0.38% 0.013
llama2-70B-diag 0-shot 31.74% 73.49% 0.058 20.79% 6.29% 0.105
Table 2: Evaluation Results of Models on MR-GSM8k: This table presents a detailed breakdown of each model’s performance, including True Positive Rate (TPR), True Negative Rate (TNR) and Matthews Correlation Coefficient. The ’Step’ column represents the accuracy of correctly identifying an incorrect solution and pinpointing the first error step. ’S+R/M’ column showcased the accuracy of not only locating the first error step in incorrect solutions but also correctly explaining the error’s rationale. The overall MR-Score/M is a normalized metric ranges from 0 to 1 and calculated based on formula described in Section-3. The M here means the error reason is labelled manually and MR-Score is calculated based on the results of that.
Figure 4: Performance Visualization on Incorrect MR-GSM8k Solutions: This graph depicts the performance of Claude2, GPT3-5, and GPT4 in identifying incorrect solutions in MR-GSM8k. ’TNR’ denotes the True Negative Rate. ’TN+Step’ refers to the ratio where models correctly identified an incorrect solution and located the first error step. ’TN+Step+Reason’ represents the frequency of models correctly determining solution correctness, identifying the first error step, and providing an accurate error reason.
Figure 5: Error Rate Analysis of Closed-Source Models on MR-GSM8k: This figure illustrates the True Positive Rate (TPR), True Negative Rate (TNR), False Positive Rate (FPR), and False Negative Rate (FNR) of closed-source models evaluated on MR-GSM8k, offering insights into their accuracy in various aspects of solution evaluation.
@@ 4.2 Open-sourced LLMs Evaluation
As mentioned briefly in Section 1 and 2, recent works of fine-tuning language models with program of thought, different reasoning types, and different difficulties have all proven effective in enhancing GSM8K benchmark results. However, just like GPT3-5 has similar pass@1 rate with GPT4 in GSM8K test set, will these open-sourced models specialized in math reasoning performs differently in the MR-GSM8k?
To address this query, we selected several state-of-the-art open-source models fine-tuned on the 70B Llama architecture (Touvron et al., 2023). Namely, they are WizardMath-70B(Luo et al., 2023), Mammonth-70B(Yue et al., 2023) and MetaMath-70B(Yu et al., 2023). WizardMath-70B underwent training on augmented problems of varying difficulty levels and was further refined through a process-oriented reinforcement learning procedure. Mammoth-70B received training on a diverse range of STEM-related problems, including those from the GSM8K dataset, which were augmented with GPT4-generated code solutions. MetaMath-70B’s training encompassed a substantial volume of augmented data, including rephrased and backward-transformed problems from the GSM8K set. Given that these models are specialized and not fine-tuned for general instructions, we employed three-shot in-context learning examples during inference to guide the models in adhering to the desired format and reasoning logic.
As demonstrated in Table-2, despite having exposure to similar training data, all three models failed miserably in the MR-GSM8k benchmark. This outcome highlights their lack of generalization capabilities when faced with problems similar to their training data but presented in different formats. Among these models, MetaMath displayed the most commendable performance, which is not entirely unexpected. Over half of its training dataset (approximately 240k instances) is derived directly from the GSM8K dataset, including a balanced mix of answer augmentations, question rephrasings, and backward transformations of questions. Despite having a training dataset with question types akin to those in MR-GSM8k (e.g., original and POT questions), the Mammoth model failed to demonstrate a deeper meta-understanding of the data, correctly answering only one out of 1573 incorrect solutions. WizardMath, trained using the Proximal Policy Optimization algorithm (Schulman et al., 2017) to optimize the joint reward of instruction adherence and solution process, also underperformed on our benchmark.
@@ 4.3 In Domain Finetuning
For our base model, we utilized Llama-2-70B-base, aligning with the approach of other open-source SOTA math models. We merged the GSM8K training set with the GPT-4 generated diagnostic data that composed of 5k incorrect solutions and 4k correct solutions. For fine-tuning, we adopted the Qlora method (Dettmers et al., 2023), maintaining the same hyperparameters as used for MetaMath-70B. The outcomes of this approach are detailed in the last row of Table-2. Remarkably, despite the zero-shot setting and a considerably smaller training set, the fine-tuned Llama2 model outperformed all open-source models and even surpassed GPT3-5. Notably, the fine-tuned Llama2 model demonstrated a tendency distinct from that of GPT3-5 and other open-source models; it was less inclined to accept solutions but to over reject solutions regardless of the correctness. As depicted in Figure-6, of the 99 questions where the model accurately predicted both correctness and the first error step, a significant portion involved questions with POT and backward reasoning types. This is particularly noteworthy given that the model was trained exclusively on original questions.
@@
Figure 6: Problem types of incorrect solutions that llama2-70B-MR models has successfully found error step and error reason. Note the training set only includes solutions from original problem.
@@
Caution is necessary when interpreting the outcomes of in-domain fine-tuning. Although the fine-tuned model achieved results comparable to GPT3-5 post-fine-tuning, it’s important to note that the overall number of correct diagnoses for incorrect solutions remains relatively low (e.g. 6.29%). This underscores the challenging nature of our MR-GSM8k benchmark, where effective diagnosis across diverse solution spaces requires a comprehensive understanding of the problem. Consequently, simple fine-tuning strategies may not yield substantial improvements in performance. Another significant observation, as indicated in Table-3, is that despite outperforming GPT3-5 in MR-GSM8k, the model’s accuracy on the GSM8K test-set still trails behind that of GPT3-5. This discrepancy could be partially attributed to a lower True Positive Rate in correct solutions and a reduced efficacy in articulating error reasons for incorrect solutions. The fine-tuning process may have enabled the model to replicate diagnostic behaviors, but this approach does not seem to enhance the model’s fundamental comprehension of mathematical reasoning. Another interpretation worth considering is that language models, to a large extent, continue to function as ’stochastic parrots’ (Bender et al., 2021), lacking genuine understanding of the content they generate. It would be intriguing to explore how scaling up the diagnostic data or employing more sophisticated training methodologies might alter these outcomes. We will delve further into this discussion in the following section.
Models GSM8k MRS/M
GPT4 92.0% 0.495
Claude2 88.0% 0.191
llama2-70B-MR 74.3% 0.105
GPT3-5 80.8% 0.097
MetaMath-70B 82.3% 0.013
Mammoth-70B 76.7% 0.012
WizardMath-70B 81.6% 0.001
@@
Table 3: Comparison of the performances of SOTA models on GSM8K and MR-GSM8k. The GSM8K results are retrieved from corresponding paper. llama2-70B-GSM8K are llama2-70B model finetuned on the training set of GSM8k only. llama2-70B-MR are llama2-70B model finetuned on the GSM8k training set and its meta-reasoning augmentation by GPT4. The MRS/M stand for the MR-Score in MR-GSM8k calculated from manual labelling results.
5 Ethical Discussion and Limitations
Ethical and Societal Dimensions of Meta-Reasoning Evaluation
As briefly discussed in Section-2, the performance on math reasoning tasks are often chosen as a key indicator of the general competence of the evaluated language model, due to its inherent requirements on cognitive abilities such as natural language understanding, concept abstraction, deductive reasoning etc. A common underlying assumption of this logic is that models succeed in reasoning tasks may possess the relevant cognitive/meta capabilities, and such capabilities should remain effective across different domains. Put aside the potential controversies behind this logic, we would like to contend that it is not sufficient to unveil the cognitive depth of the evaluated model by only looking at the computation results. What becomes equally important is the validity and logic of the reasoning process employed by the evaluated model. Therefore, the "reason about reasoning" paradigm emerges as a vital meta-evaluative tool. It allows us to delve deeper into the reasoning process itself, probing the model’s ability to apply learned concepts flexibly and logically across different scenarios.
As illustrated by Figure-2, for the evaluated model to successfully diagnose the solution correctness, it is necessary for the model to be able to infer the correct result and also be able to counterfactually reasons along different reasoning paths and actively examines the conditions and assumptions made on different steps. It is unlikely to succeed in this paradigm without a holistic understanding and robust mastery of the underlying concepts.
Moreover, the meta-reasoning paradigm also opens up a window for the researchers to examine the "alignment" and "grounding" levels of the evaluated model. In subjects of natural science, we would expect an intelligent agent that aligns epistemically with humans to ground its logic on the rules and principles that closely matches the physical world we live in. A quick example would be the unit inconsistency issue during calculations of models (Toh et al., 2023) which highlights a lack of fundamental ontological understanding. In the downstream application scenarios such as education and consultancy, such limitations might present a fundamental deficiency that prevents a larger adoption of LLM technologies.
Limitations of the ’Reason About Reasoning’ Evaluation Paradigm and MR-GSM8k Dataset